OpenStudy (anonymous):

help me out... It is given that -1 is one of the roots of quadratic equation mx^2 - 3x - k = 0 where m and k are constants. Express m in term of k.

hartnn (hartnn):

just plug in x=-1 in that equation as -1 is the root, it satisfies that equation

OpenStudy (anonymous):

is it the answer ... m+3-k=0

hartnn (hartnn):

yes, so m=k-3

OpenStudy (anonymous):

ok, how about this question.....  express y in terms of x. how do i do this?

hartnn (hartnn):

tricky.... let xy = Y 1/x=X you have 2 points, in XY plane, find the equation of line in terms of X and Y, can you ?

OpenStudy (anonymous):

no

hartnn (hartnn):

you don't know how to find equation of line when 2 points are given ? start by finding the slope, have the formula of slope with you ?

OpenStudy (anonymous):

is the slope of the line is the gradient ? so can i make it using y-y=m(x-x)

hartnn (hartnn):

yes, correct, use that and tell me the equation of line...

OpenStudy (anonymous):

y-3= 5(x-2) y-3=5x-10 y=5x-7

hartnn (hartnn):

correct! its actually Y=5X-7 now plug in back, X = 1/x , Y =xy

OpenStudy (anonymous):

xy= 5/x - 7

hartnn (hartnn):

yes, just isolate the 'y' by dividing both sides by 'x'

OpenStudy (anonymous):

5/x^2 - 7/x

hartnn (hartnn):

yes, i think thats correct.